THEORETICAL STUDY OF LOW-LYING STATES OF UH

\underline{F. MICHAEL DOLG} and XIAOYAN CAO, Institut für Theoretische Chemie, Universität zu Köln, Greinstr. 4, D-50939 Köln, Germany..

Large-scale multireference configuration-interaction (MRCI) calculations using the all-electron scalar-relativistic Douglas-Kroll-Hess (DKH) Hamiltonian, as well as a relativistic energy-consistent small-core pseudopotential (SPP) for uranium, have been performed to study the low-lying electronic ΛS states of uranium monohydride UH with term energies below 0.5 eV. After taking spin-orbit effects into account both DKH/MRCI and SPP/MRCI calculations predict a ^{4}I_{9/2} ground state. The calculated ground state molecular constants of both approaches show a good agreement with each other (MRCI+Q, DKH: Re=2.021 Å, ω_{e}=1483 cm^{−1}, D_{e}=2.79 eV; SPP: R_{e}=2.011 Å, ω_{e}=1497 cm^{−1}, D_{e}=2.85 eV), as well as with available experimental data (ω_{e}=1424 cm^{−1} in argon matrix).